OpenStudy (anonymous):

the perimeter of a rectangle is 26ft. the length is 9 ft longer than the width. find the dimensions. Write a system of linear equations and solve the resulting system. Let x be the length and y be the width. write the first equation. 2x+2y= write the second equation x=y +_____. what is the length? ___ ft what is the width? ____ft

OpenStudy (anonymous):

Let's let P rep. the perimeter, l rep. the length, and w rep. the width. We know from the question that P = 26 We also know that because it is a rectangle: P = 2*l + 2*w Therefore 26 = 2*l + 2*w The second equation tells us that the length is 9ft longer than the width: Therefore l = w + 9 Does that help?

OpenStudy (anonymous):

I didn't read the question: replace x with l and y with w

OpenStudy (anonymous):

do you know what 2x+2y= and x=y+____, what is the length? ___ft what is the width? ___ft

OpenStudy (anonymous):

I was hoping you could tell me. Since you have two equations and two "unknowns" you can solve for them. Try isolating for w or l and substituting that into the other equation.

OpenStudy (anonymous):

i'm lost on this question

OpenStudy (anonymous):

did you understand my first hints? "Let's let P rep. the perimeter, l rep. the length, and w rep. the width. We know from the question that P = 26 We also know that because it is a rectangle: P = 2*l + 2*w Therefore 26 = 2*l + 2*w The second equation tells us that the length is 9ft longer than the width: Therefore l = w + 9" ???

OpenStudy (anonymous):

no, I am horrible at math

OpenStudy (anonymous):

Ok, do you understand what it means to let a variable represent a number. In this case a measurement in feet ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

a variable is just a letter we use in math to represent a number that we don't necessarily know

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so we are trying to find a length and a width so I'll call them L and W

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(and P will represent the perimeter)

OpenStudy (anonymous):

Perimeter is the total distance around something

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Rectangles have two longer sides(lengths) and two shorter sides(widths) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So P = L + L + W + W right?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

this is the same as saying: P = 2L + 2W right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

so because the perimeter is given in the equation we can change this from: P = 2L + 2W to: 26 = 2L + 2W

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

excellent. that is 1 of the 2 equations. Now the next one: since the length(L) is 9 feet longer than the width(W) we know that the width + 9 feet is the length do you follow?

OpenStudy (anonymous):

but how do we get to the answer 2x+2y=___?

OpenStudy (anonymous):

that we got from the first equation. Since x and y and length and width: 26 = 2L + 2W is the same as 2L + 2W = 26 or 2X + 2Y = 26

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

are we working on how to get x=y +___?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if the length and width were the same then this would be true right? L = W

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you still there?

OpenStudy (anonymous):

yes, sorry. now since the length is 9 feet longer than the width we can say that the width plus 9 is the length. right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

great, this statement: "the width plus 9 is the length" is the same as the equation: L = W + 9

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now you have both equations: 26 = 2W + 2L and L = W + 9

OpenStudy (anonymous):

so is it x=y+9?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so how do we now get the length and width?

OpenStudy (anonymous):

you still there?

OpenStudy (anonymous):

yes, sorry answering two questions at once. Now, isolate for one of the variables in one of the equations: L = W + 9 (this is already isolated for L, yey!, don't have to do anything) let's put this into the other equation: everywhere I see L I'll replace it with W + 9

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so doing that: 2W + 2L = 26 2W + 2(W + 9) = 26

OpenStudy (anonymous):

so the width = 26?

OpenStudy (anonymous):

or 35?

OpenStudy (anonymous):

2W + 2(W + 9) = 26 distribute: 2W + 2W + 18 = 26 subtract 18 from both sides: 2W + 2W + 18 -18 = 26 - 18 on the left side the 18s cancel: 2W + 2W = 26 - 18 on the right side we subtract: 2W + 2W = 8 collect the w's: 4W = 8 divide both sides by 4: (4/4)W = 8/4 simplify: W = 2

OpenStudy (anonymous):

so width = 2 ft?

OpenStudy (anonymous):

sttill there, I only have 5 more minutes on the computer to finish this question?

OpenStudy (anonymous):

yes 2ft. now solve for L: 26 = 2L + 2W 26 = 2L + 2(2) 26 = 2L + 4 26 -4 = 2L 22 = 2L L = 11

OpenStudy (anonymous):

so length is 11 ft?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, i really appreciate your help. thank you very much

OpenStudy (anonymous):

no problem. you'll get the hang of it